OpenStudy (anonymous):

if the diameter of a circle has endpoints A (7,2) and B (-1,8), where is the center

OpenStudy (anonymous):

What's the average of 7 and -1 (x coordinate) What's the average of 2 and 8 (y coordinate)

OpenStudy (whpalmer4):

You need three formulas: Distance formula between two points is given by \[d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Midpoint of a line: \[(x,y) = (\frac{1}{2}(x_1+x_2), \frac{1}{2}(y_1+y_2))\] Equation of a circle: \[(x-h)^2+(y-k)^2=r^2\]where center is at \((h,k)\) and radius \(r\) Use the midpoint formula to find the center of the circle. Use distance formula to find the radius (distance between center and one endpoint of a diameter). Use the circle formula to write the formula of the circle.

OpenStudy (whpalmer4):

Smart-aleck answer: "in the middle" :-) Looks like you don't have to do all of the work I equipped you to do, but someday soon you will :-)

OpenStudy (anonymous):

use the mid point formula that says \[x=\frac{ X_{1} +X _{2}}{ 2 }, y=\frac{ Y_{1}+Y _{2} }{ 2 }\]

OpenStudy (anonymous):

i dont know which ones x or y

OpenStudy (anonymous):

 you gotta find the mid point if you got ends of the circle

OpenStudy (anonymous):

x,y are the coordinates of mid point that you have to find

OpenStudy (anonymous):

and x1, y1 are (7,2) and x2,y2 are (-1,8) substitute the values and find the answer

OpenStudy (anonymous):

x co ordinate of your mid point will be (7+(-1))/2=6/2=3 similarly find y

OpenStudy (anonymous):

@kbarone I already said you can do 2 simple averages to find your answer: What's the average of 7 and -1 (x coordinate) What's the average of 2 and 8 (y coordinate) (x,y) will be the answer

OpenStudy (anonymous):

(B)? (3,5)?